OpenStudy (anonymous):

Mary walked north from her home to Sheila's home, which is 4.0 kilometers away. Then she turned right and walked another 3.0 kilometers to the supermarket, which is 5.0 kilometers from her own home. She walked the total distance in 1.5 hours. What were her average speed and average velocity?

OpenStudy (anonymous):

 Average speed = total distance / total time = 3+4/1.5= 5.666667 Km per hour. Average velocity= displacement / total time = 5/1.5 = 3.3333 Km per hour. It is so easy can't you really solve it ??